OpenStudy (anonymous):

Write the equation of the circle in standard form.  Find the center, radius, intercepts, and graph the circle. x^2 + y^2 + 10x +8y + 16 = 0

OpenStudy (anonymous):

1. Move the 16 to the right side of the equation. 2. Group terms with similar variable. 3. Then find the perfect square of each variable (b/2)^2, for x-terms that would be (10/2)^2=25, for the y-term it would be (8/2)^2 = 16. 4. Add 16 and 25 to both sides. 5. Find the perfect squares for each variable (e.g x^2+10y+25 = (x+5)^2). What is y^2+8y+16 as a perfect square? 6. Then you should get the answer.

OpenStudy (anonymous):

Are you working on the question using the steps above? at what point are you lost?

OpenStudy (anonymous):

I don't understand the way how your explaining it.

OpenStudy (anonymous):

you lost me at step 3

OpenStudy (anonymous):

Step 1: x^2+10x+y^2+8y+16=0 Step2: x^2+10x+y^2+8y=-16 (move the 16 to the right hand side by subtraction) Step3: For the x-term 'b' =10 For the y-term 'b' = 8 So to find the number that would make the x-term and y-term perfect squares we use (b/2)^2 So we get: x^2+10x+(10/2)^2+y^2+8y+(8/2)^2 = -16+(10/2)^2+(8/2)^2 (x^2+10x+25)+(y^2+8y+16) = -16+16+25 (x^2+10x+25)+(y^2+8y+16) = 25 Do you understand what I did here?

OpenStudy (anonymous):

From here you just need to find the perfect squares for both terms in the parentheses and that should be your answer.

OpenStudy (anonymous):

yes I do now thanks

OpenStudy (anonymous):

Sure no problem